# I’m Sorry Dave, I’m Afraid I Can’t Return That: On YouTube Search API Use in Research

### v1

@@ Abstract
YouTube is among the most widely-used platforms worldwide, and has seen a lot of recent academic attention. Despite its popularity and the number of studies conducted on it, much less is understood about the way in which YouTube’s Data API, and especially the Search endpoint, operates. In this paper, we analyze the API’s behavior by running identical queries across a period of 12 weeks. Our findings suggest that the search endpoint returns highly inconsistent results between queries in ways that are not officially documented. Specifically, the API seems to randomize returned videos based on the relative popularity of the respective topic during the query period, making it nearly impossible to obtain representative historical video samples, especially during non-peak topical periods. Our results also suggest that the API may prioritize shorter, more popular videos, although the role of channel popularity is not as clear. We conclude with suggested strategies for researchers using the API for data collection, as well as future research directions on expanding the API’s use-cases.
1 Introduction
YouTube has been used to study a wide array of crucial societal problems like online hate [1, 17], accessibility [13], pseudoscientific misinformation [18], online scams [10], and child exposure to inappropriate content [5, 9, 16]. Much of this work has made use of the YouTube Data API, which offers several endpoints.
@@ 1 Introduction
The way we understand these problems is influenced by the quality of the data we can get from the platform, yet researchers have pointed out that the YouTube API can behave inconsistently and return seemingly biased results [3, 24]. Beyond documentation and implementation errors [12], some works suggest that some of these biases may result from systematic API behavior, particularly through the keyword-based Search: list endpoint (“search” henceforth), that makes obtaining random samples very difficult [14, 22].
@@
Although YouTube data are collected in a variety of ways, including crawling through recommended videos [11, 21, 23], deploying sockpuppet accounts [7, 8], extensions installed on participants’ browsers [6], or collecting videos from pre-curated channels [2, 9], the search endpoint remains an important part of the data collection pipeline for a lot of research [1, 13, 15, 16, 17, 19, 24]. Given its importance and the fact that its behavior is poorly understood [3, 14], we conduct an audit of this API endpoint. We run identical queries at 5-day intervals across a period of approximately 3 months to answer the following research questions:
@@
    How consistent and reliable are the data returned by the Search endpoint?
@@
2 Background & Related Work
Based on the official documentation, [^1] https://developers.google.com/youtube/v3/docs/search/list the search endpoint allows a user to search by keywords, location, or live events, and enables filtering by several other parameters such as date ranges, specific channels, etc. However, this endpoint has a quota cost of 100 units per query; this is considerably higher than ID-based endpoints, which typically only cost 1 quota unit per query. With the default daily quota for a newly created client being 10,000, this only allows 100 search queries per day per client. However, the YouTube Data API has a researcher access program that provides substantially higher quotas to vetted accounts. This endpoint is not designed for volume, and the maximum number of results per query is 500 (a maximum of 50 per page and a maximum of 10 pages) [25].
One approach for comprehensive data collection was to identify topical “seed” channels and videos (either through external sources [16] or keyword search [18]), and obtain videos recommended by YouTube as being relevant to those initial sets to expand the dataset. However, YouTube deprecated the relatedToVideoId response that enabled this approach in 2023, effectively eliminating it from being conducted through the API.
An alternative advocated strategy has been time-split queries for clients endowed with sufficient quota [25]. The API allows the client to add parameters for publishedAfter and publishedBefore to restrict the data collection period. Researchers have used this to query in a “one per X time” fashion [22], where the observation period is split into time bins, each of which is queried separately to circumvent the 500-video limit [20, 24]. In theory, this should enable researchers to obtain every video uploaded on the topic, unless more than 500 videos were uploaded on a specific day.
Another commonly used strategy is identifying relevant channels through external sources like SocialBlade [9] or Reddit [2] and querying the API for their videos. While this can be done using several endpoints, for example, using Channels: list to extract a “playlist” of a channel’s uploaded videos and then querying the PlaylistItems: list endpoint for those videos, or adding a “channelId” parameter to the search endpoint, few papers clarify the exact endpoints used [13] or the dates on which the queries themselves are made [18]. As we later show, both of these can influence returned data.
3 Methods
For our experiments, we choose a diverse range of political, scientific, and entertainment topics that are either regional or international and vary in size and recency:
  • Black Lives Matter (BLM)
  • Brexit
  • US Capitol Riots
@@ 3 Methods
For each topic, we run the same query every five days through the YouTube Data API v3 Search: list endpoint, starting on February 9 and ending on April 30, 2025 (see Appendix A for exact queries). Due to a technical problem, the collection on April 5th was skipped. In each query, we obtain a focal date (e.g., the day of the referendum for Brexit), and we set our data collection period between two weeks before and after this focal date (i.e., a total collection span of 28 days per topic). We then send queries for every hour within these 28 days to circumvent the maximum limits in returned videos imposed by the API following the strategy outlined in Section 2, resulting in 4032 total queries for every collection (24 hours $\times$ 28 days $\times$ 6 topics). This results in 16 snapshots over a period of 12 weeks. All queries are set to return results in reverse chronological order. We choose this order as it is an immutable video property, whereas other ordering options like view count or relevance may change over time. Thus, it offers the best baseline to study API consistency.
We then compare the consistency in the videos returned per collection using video ID set similarities. We find that this similarity decays over time, indicating that datasets collected using the exact same historical query may differ vastly based simply on when the queries were made. Moreover, we analyze the historical time points on which the most data is returned, finding that the API may be suppressing data returns for periods of relative topical inactivity even though returning these data would not bring the response above the maximum number of videos allowed by the API. Using a second-order Markov chain to model “transitions” between the presence or the absence of a video in successive collections, we find that video omission or inclusion is mostly conducted in a “rolling window” fashion. Finally, we analyze whether any video metadata (e.g., likes, views, etc.) is associated with more consistent video returns, finding that the API is more likely to return more popular videos that are drawn from less active topics.
4 API Behavior
In this section, we highlight how the API’s search endpoint returns inconsistent results. We document some of the potential mechanisms of this inconsistency and how videos may be drawn when determining what to return to the client.
4.1 Temporal Inconsistencies
For every collection instance at time $t$, we obtain the set of video IDs returned $S_{t}$ and calculate its Jaccard similarity with the set obtained in the previous collection $S_{t-1}$ and the very first collection $S_{t-n}$. We plot these rolling Jaccard similarities in Figure 1. In addition, we show descriptive statistics for the number of videos returned in Table 1.
Figure 1: Jaccard similarities of video ID sets relative to the previous and the first collection instance using the “search” endpoint. “Error bars” represent the set difference of $S_{t-1}-S_{t}$ (bottom bars) and $S_{t}-S_{t-1}$ (top bars).
@@ 4.1 Temporal Inconsistencies
Table 1: Descriptive statistics for number of videos returned per topic across collections.
We find some non-constant differences between successive runs. However, over time, these differences compound to form video sets that are vastly different from the initial collection, to Jaccard values as low as $\sim$0.3 after 3 months. This equates to only 46% of the videos per set being shared, with the majority being unshared. The exception is the Higgs topic, which retains much higher consistency than the rest–we offer an explanation of why that may be in Section 5. As shown by the “error bars” in Figure 1, content deletions cannot explain this phenomenon as we find videos at $t$ that were not seen at $t-x$, despite these queries being fully historical (i.e., not spanning the query date). Moreover, we verify that this behavior does not extend to other endpoints that use ID-based queries (see Appendix B).
4.2 Randomization Mechanisms
An obvious question is whether the discrepancies are due to ceiling effects, despite the query specifying that results should be returned in chronological order. That is, if the eligible videos for a particular query exceed the maximum of 50 per page/500 per query allowed by the API, then we might expect that the results returned will be randomized. In that case, we should expect that some of our queries return more than 500 eligible videos, and that Jaccard similarities between collections are lower for hours on which more videos are uploaded. In this section, we test this theory.
We first obtain some descriptive statistics of videos returned for each hour and each of the topics (Table 2). The maximum number of videos returned for any given hour (max) remains well below the theoretical maximum of 50 per page, ruling out the ceiling effect explanation. Moreover, we compute the Spearman coefficient $\rho$ for the correlation between the Jaccard similarity of sets $T_{1}$ (first collection) and $T_{L}$ (last collection) and the average number of videos returned for that hour, after dropping all hours for which 0 videos are returned across collections to avoid inflating Jaccard similarity values based on empty sets. Based on the above reasoning, we should expect negative correlations between these variables. However, we observe almost the exact opposite pattern: For all but the Higgs topic, for which we observe a non-significant (negative) correlation, there are weak positive correlations between the number of videos for that hour and the Jaccard similarity, indicating that similarity values are, on average, higher for busier hours. Although this may be an artifact of the idea that more videos simply stabilize the Jaccard value, this analysis demonstrates that fewer videos do not necessarily mean that those videos will be the same across collections.
@@ 4.2 Randomization Mechanisms
Table 2: Descriptive statistics for per-hour number of videos returned. *p ¡ 0.05, **p ¡ 0.01, ***p ¡ 0.001. N is the number of videos retained after all hours with no videos returned across collections are dropped.
We also plot the daily video frequencies of the first and last collections alongside the Jaccard similarity for these daily sets in Figure 2. To ensure that these are not edge cases, we further plot the average daily frequencies across all collections. As can be seen, the average daily frequency distributions per collection map almost perfectly on each other. However, the volume of videos returned does not map onto the Jaccard similarities in any consistent manner, confirming our above findings of weak or non-existent correlations between these two factors. Most videos returned are uploaded around the focal date, with the exception of the BLM topic (likely due to when the protests surrounding George Floyd’s death intensified; the topical peak is recorded on Blackout Tuesday). Jaccard similarities at frequency peaks are comparable to other days on which much fewer videos are uploaded, indicating that the YouTube Data API operates on time-dependent systematic randomization. Interestingly, the YouTube recommendation system uses an empirical distribution of video popularity against its age to determine how relevant it is to recommend to users [4]. These patterns suggest that the API may similarly sample from topic-wide empirical distributions in terms of over-time interest.
Figure 2: Daily frequencies of videos returned with daily Jaccard similarities between first and last collections. Dashed vertical lines represent the corresponding topic’s D-day.
Overall, we do not find evidence of ceiling effects driving the API discrepancies. Instead, our results suggest that the YouTube API samples videos from empirical distributions, returning results based on the relative density of topical interest and even forcing zero videos to be returned when this relative density is adequately low. However, it is unclear how this interest is computed (e.g., if it is the volume of videos uploaded or something else).
4.3 Attrition Analysis
@@
Figure 3: Transition probabilities for presence or absence of videos in a second-order Markov chain. Columns show outgoing probabilities to the rows.
@@
5 Factors Behind Return Likelihood
@@
Our next analysis focuses on whether the YouTube API is more likely to return videos with certain features at higher rates. For each video per topic, we count the number of times it is returned in our API calls, and we use this frequency as our dependent variable. Frequency ranges from 1-16 (contingent on our number of collections), and we split it into four roughly equal bins (1-5, 6-10, 11-15, 16), taking into account that 16 is the modal value . We then perform an ordinal regression using a logit link function (hence the binning). For robustness, we also perform an unbinned ordinal regression and a multiple regression taking frequency as a continuous variable; we report these models in Appendix C and find largely consistent results. As candidate predictors, we use the video-level features of topic, video duration, video definition, number of views, comments, and likes; and the channel-level features of channel views, subscribers, and number of videos uploaded. All continuous features are log-transformed to reduce multicollinearity and standardized for better comparison between coefficients.
Variable $\beta$ SE 95% CI
SD (quality) -0.018 0.079 [-0.173, 0.137]
brexit (topic) ***1.231 0.098 [1.039, 1.423]
capriot (topic) -0.160 0.093 [-0.341, 0.022]
grammys (topic) *0.171 0.083 [0.008, 0.333]
higgs (topic) ***3.10 0.141 [2.826, 3.379]
worldcup (topic) 0.161 0.101 [-0.037, 0.359]
duration ***-0.115 0.028 [-0.170, -0.061]
views 0.161 0.088 [-0.011, 0.333]
likes **0.285 0.095 [0.098, 0.471]
comments 0.069 0.064 [-0.058, 0.195]
channel age 0.049 0.031 [-0.012, 0.110]
channel views *0.3176 0.135 [0.053, 0.582]
channel subs **-0.3784 0.122 [-0.617, -0.140]
# channel videos -0.0212 0.075 [-0.169, 0.126]
Table 3: Standardized regression coefficients for binned ordinal model. *$p<0.05$, **$p<0.01$, ***$p<0.001$.
Starting with video features, we see a negative effect of video duration and a positive effect of number of likes, indicating that shorter, more widely-liked videos are returned in more collections. We should note that the non-significant effect of views and comments is likely due to their very high correlations and thus shared variance with the number of likes ($r=0.92,p<0.001$ and $r=0.89,p<0.001$, respectively). Indeed, these predictors become significant when likes are dropped from the model, suggesting that more popular videos overall are returned more often. For channel features, the total number of channel views also has a positive effect, while the number of channel subscribers has a negative one. Curiously, despite a near-perfect correlation between these two variables ($r=0.97,p<0.001$), the negative effect of channel subscriptions persists when total channel views are dropped; however, total channel views become non-significant when subscriptions are dropped. For this reason, we urge caution in interpreting channel-related results as they may be spurious.
@@
With respect to topics, we find that the Brexit, Grammy Awards, and Higgs Boson topics are returned in significantly more collections than BLM. Capitol Riot and World Cup show no effects. Although these topics look unrelated at first, a deeper look reveals that they are linked by one factor: size. We determine this using metadata returned with every query we send. As part of this metadata, the YouTube API returns a value called pageInfo.totalResults which reflects “the total number of results in the result set”, [^2] https://developers.google.com/youtube/v3/docs/search/list i.e., the total number of videos that match our query, with a maximum potential value of 1,000,000. We obtain the minimum, maximum, mean, and modal values from this field across every hour and collection run that we query per topic. As seen in Table 4, the three topics whose videos have higher appearance frequencies are also the smallest (and the only ones without a modal value of 1M, which is the maximum). Although this observation requires further experiments to verify, the results indicate that queries with a smaller video pool to draw from may return more consistent results (and may explain why Higgs, which is smaller by orders of magnitude compared to other topics, is also by far the most consistent in Figure 1).
@@
Table 4: Potential video pool size per topic.
@@
It is also interesting to note that, despite the modal number of videos returned across all collection hours for all topics being 0, the modal value for the available pool is much higher (and the maximum of 1M for 3/6 topics). Since it is unlikely that 1M videos are consistently uploaded every hour for any of these topics, this behavior suggests that the API does not take into account time constraints in determining the total pool of available videos.
@@
Comparing Tables 1 and  4, it is also striking that the number of videos returned is much closer across topics than their respective topic sizes suggest, which is consistent with our distribution density explanation. This becomes even more apparent when scrutinizing the y-axis on Figure 2: The most-populated peaks are recorded for topics where the rest of the time-series is relatively inactive (e.g., Capitol Riot, Grammys), while topics that are more active throughout (e.g., World Cup, which is an ongoing tournament rather than a one-off event) record peaks at lower absolute values, most likely due to the number of videos to be returned being fixed and the actual videos being drawn from an empirical distribution.
6 Discussion

### v2

# On YouTube Search API Use in Research

@@ Abstract.
YouTube is among the most widely-used platforms worldwide, and has seen a lot of recent academic attention. Despite its popularity and the number of studies conducted on it, much less is understood about the way in which YouTube’s Data API, and especially the Search endpoint, operates. In this paper, we analyze the API’s behavior by running identical queries across a period of 12 weeks. Our findings show that the search endpoint returns highly variable results between queries. Specifically, the API seems to randomize returned videos based on the relative popularity of the respective topic during the query period, making it nearly impossible to obtain representative historical video samples, especially during non-peak topical periods. Our results also suggest that the API may prioritize shorter, more popular videos, although the role of channel popularity is not as clear. We conclude with suggested strategies for researchers using the API for data collection, as well as future research directions on expanding the API’s use-cases.
YouTube, API, Research tools, Data collection
journalyear: 2025copyright: ccconference: Proceedings of the 2025 ACM Internet Measurement Conference; October 28–31, 2025; Madison, WI, USAbooktitle: Proceedings of the 2025 ACM Internet Measurement Conference (IMC ’25), October 28–31, 2025, Madison, WI, USAdoi: 10.1145/3730567.3764492isbn: 979-8-4007-1860-1/2025/10ccs: Information systems Information retrievalccs: Information systems World Wide Webccs: Human-centered computing HCI design and evaluation methodsccs: Human-centered computing Collaborative and social computing
1. Introduction
@@
YouTube has been used to study a wide array of crucial societal problems like online hate (Albadi et al., 2022; Papadamou et al., 2021), accessibility (May et al., 2024), pseudoscientific misinformation (Papadamou et al., 2022), online scams (Liu et al., 2024), and child exposure to inappropriate content (Gkolemi et al., 2022; Khan et al., 2024; Papadamou et al., 2020). Much of this work has made use of the YouTube Data API, which offers several endpoints. However, researchers have pointed out that the YouTube API may sometimes return seemingly biased results (Choi et al., 2025; Violot et al., 2024). Beyond documentation and implementation errors (Martin-Lopez et al., 2021), some works suggest that this can result from systematic API behavior, particularly through the keyword-based Search: list endpoint (“search” henceforth), which makes obtaining random samples difficult (McGrady et al., 2023; Rieder et al., 2018).
@@
Although YouTube data are collected in a variety of ways, including crawling through recommended videos (Mamié et al., 2021; Ribeiro et al., 2020; Röchert et al., 2020), deploying sockpuppet accounts (Hussein et al., 2020; Jung et al., 2025), extensions installed on participants’ browsers (Hu et al., 2024), or collecting videos from pre-curated channels (Bertaglia et al., 2024; Khan et al., 2024), the search endpoint remains an important part of the data collection pipeline for a lot of research (Albadi et al., 2022; May et al., 2024; Nigatu and Raji, 2024; Papadamou et al., 2020, 2021; Pera and Aiello, 2024; Violot et al., 2024). Given its importance and the fact that its behavior is poorly understood (Choi et al., 2025; McGrady et al., 2023), we conduct an audit of this API endpoint. We run identical queries at 5-day intervals across a period of approximately 3 months to answer the following research questions:
@@
    How consistent are the data returned by the Search endpoint?
@@
Comparing the set similarities of videos obtained with each query, we find that this similarity decays over time, indicating that datasets collected using the exact same historical query may differ vastly based simply on when the queries were made. Moreover, we analyze the historical time points on which the most data is returned. We find that the API may be withholding data for periods of relative topical inactivity even though returning these data would not bring the response above the maximum number of videos allowed by the API. Using a second-order Markov chain to model “transitions” between the presence or the absence of a video in successive collections, we find that video omission or inclusion is mostly conducted in a “rolling window” fashion. Finally, we analyze whether any video metadata (e.g., likes, views, etc.) are associated with more consistent video returns, finding that the API is more likely to return more popular videos that are drawn from less active topics.
@@
2. Background
Based on the official documentation, [^1] https://developers.google.com/youtube/v3/docs/search/list the search endpoint allows a user to search by keywords, location, or live events, and enables filtering by several other parameters such as date ranges, specific channels, etc. However, this endpoint has a quota cost of 100 units per query; this is considerably higher than ID-based endpoints, which typically only cost 1 unit. With the default daily quota being 10,000, this allows 100 search queries per day per client. However, the YouTube Data API has a researcher access program that allows higher quotas to vetted accounts. This endpoint is not designed for volume. The maximum number of results per query is 500 (max. 50 per page and max. 10 pages) (Yin and Brown, 2018), thus, it is necessary for YouTube to sample the videos it returns.
One approach to collect all videos on a topic was to identify topical “seed” channels and videos (either through external sources (Papadamou et al., 2020) or keyword search (Papadamou et al., 2022)), and obtain videos recommended by YouTube as being relevant to those initial sets to expand the dataset. However, the relatedToVideoId response that enabled this approach was deprecated in 2023, eliminating it from being conducted through the API.
An alternative advocated strategy has been time-split queries for clients endowed with sufficient quota (Yin and Brown, 2018). The API allows the client to add parameters for publishedAfter and publishedBefore to restrict the data collection period. Researchers have used this to query in a “one per X time” fashion (Rieder et al., 2018), where the observation period is split into time bins, each of which is queried separately to circumvent the 500-video limit (Porreca et al., 2020; Violot et al., 2024). In theory, this should enable researchers to obtain every video uploaded on the topic, unless more than 500 videos were uploaded on a specific day. However, recent work finds a strong recency bias with this approach, with a much higher volume of videos for dates closer to the query date than for historical dates (Rieder et al., 2025).
Another commonly used strategy is identifying relevant channels through external sources like SocialBlade (Khan et al., 2024) or Reddit (Bertaglia et al., 2024) and querying the API for their videos. While this can be done using several endpoints, for example, using Channels: list to extract a “playlist” of a channel’s uploaded videos and then querying the PlaylistItems: list endpoint for those videos, or adding a “channelId” parameter to the search endpoint, few papers clarify the exact endpoints used (May et al., 2024) or the dates on which the queries themselves are made (Papadamou et al., 2022). As we later show, both of these can influence returned data.
3. Methods
For our experiments, we choose a diverse range of political, scientific, and entertainment topics that are either regional or international and vary in size and recency. Although these topics are not exhaustive, they allow us to observe whether certain patterns may be due to characteristics like topic sensitivity. The exact queries per topic are shown in Appendix B:
  • Black Lives Matter (BLM) (2020)
  • Brexit (2016)
  • US Capitol Riots (2021)
@@ 3. Methods
Each topic has a focal “D-day” on which a central event took place (e.g., the day of the referendum for Brexit; see Appendix B). We set our data collection period between two weeks before and after this date (i.e., a total collection span of 28 days per topic). For each topic, we run the same query every five days through the YouTube Data API v3 Search: list endpoint, starting on February 9 and ending on April 30, 2025. Due to a technical problem, the collection on April 5th was skipped. Our data and code are available on GitHub. [^2] https://github.com/alefstrat/youtube_api_audit
We send queries for every hour within these 28 days to circumvent the maximum limits in returned videos imposed by the API following the strategy outlined in Section 2, resulting in 4032 total queries for every collection (24 hours $\times$ 28 days $\times$ 6 topics) and 16 snapshots over 12 weeks. Queries are made with an API token, not OAuth 2.0, so that account effects do not confound our results. Responses are set to be returned in reverse chronological order. We choose this order as it is an immutable video property, whereas other ordering options like view count or relevance may change over time. Thus, it offers the best baseline to study API consistency. However, we stress that the YouTube API’s documentation makes no commitment to order results beyond a daily granularity. Our choice to make queries for each individual hour is so that we maintain more consistency between low-activity and high-activity days, i.e., avoiding ceiling effects for days on which more than 500 videos may have been posted. However, on such days, it can alternatively be the case that order may take precedence over time-filtering. That is, returned videos may be sampled from a given day, rather than a given hour, if the number of eligible results exceeds the maximum.
4. API Behavior
In this section, we highlight how the API’s search endpoint varies returned results. We document some of the potential mechanisms of this variability and how videos may be drawn when determining what to return to the client.
4.1. Temporal Variability
For every collection instance at time $t$, we obtain the set of video IDs returned $S_{t}$ and calculate its Jaccard similarity with the set obtained in the previous collection $S_{t-1}$ and the very first collection $S_{t-n}$. We plot these rolling Jaccard similarities in Figure 1. Table 1 shows descriptives for the total number of videos returned per collection.
Figure 1. Jaccard similarities of video ID sets relative to the previous and the first collection instance using the “search” endpoint. “Error bars” represent the set difference of $S_{t-1}-S_{t}$ (bottom bars) and $S_{t}-S_{t-1}$ (top bars).
@@ 4.1. Temporal Variability
Table 1. Descriptive statistics for number of videos returned per topic across collections.
We find differences between successive runs, which, over time, compound to form video sets that are vastly different from the initial collection. The exception is the Higgs topic, which retains much higher consistency than the rest–we offer an explanation of why that may be in Section 5. The “error bars” in Figure 1 rule out content deletions as a potential explanation; we find videos at $t$ that were not seen at txt-x, despite these queries being fully historical (i.e., not spanning the query date). This behavior does not extend to endpoints that take IDs as queries (see Appendix C).
4.2. Randomization Mechanisms
An obvious question is whether the variability is due to ceiling effects, i.e., whether randomization occurs due to the possible matches exceeding the maximum number of results allowed to be returned. In this section, we test this theory.
We first obtain some descriptive statistics of videos returned for each hour and each of the topics (Table 2). The maximum number of videos returned for any given hour (max) remains well below the theoretical maximum of 50 per page, ruling out the ceiling effect explanation. Moreover, we compute the Spearman coefficient $\rho$ for the correlation between the Jaccard similarity of sets $T_{1}$ (first collection) and $T_{L}$ (last collection) and the average number of videos returned for that hour. However, to avoid inflating Jaccard similarity values based on empty sets, we first drop all hours for which 0 videos are returned across collections. The correlation is meant as a soft test of the ceiling effect, since, if randomization is indeed more prominent when more videos are returned, we should expect lower Jaccard similarities with a higher number of returned videos (i.e., a negative correlation). However, we observe almost the exact opposite pattern: For all but the Higgs topic, for which we observe a non-significant (negative) correlation, there are weak positive correlations between the number of videos for that hour and the Jaccard similarity, indicating that similarity values are, on average, higher for busier hours. Although this may be an artifact of the idea that more videos simply stabilize the Jaccard value, this analysis demonstrates that fewer videos do not necessarily mean that those videos will be the same across collections.
@@ 4.2. Randomization Mechanisms
Table 2. Descriptive statistics for per-hour number of videos returned. *p ¡ 0.05, **p ¡ 0.01, ***p ¡ 0.001. N is the number of videos retained after all hours with no videos returned across collections are dropped.
Figure 2. Daily frequencies of videos returned with daily Jaccard similarities between first and last collections. Dashed vertical lines represent the corresponding topic’s D-day.
We also plot the daily video frequencies of the first and last collections alongside the Jaccard similarity for these daily sets in Figure 2. To ensure that these are not edge cases, we further plot the average daily frequencies across all collections. As can be seen, the average daily frequency distributions per collection map almost perfectly on each other. However, the volume of videos returned does not map onto the Jaccard similarities in any consistent manner, confirming our above findings of weak or non-existent correlations between these two factors. Most videos returned are uploaded around the focal date, with the exception of the BLM topic (likely due to when the protests surrounding George Floyd’s death intensified; the topical peak is recorded on Blackout Tuesday). Jaccard similarities at frequency peaks are comparable to other days on which much fewer videos are uploaded, indicating that the YouTube Data API operates on time-dependent systematic randomization. Interestingly, the YouTube recommendation system uses an empirical distribution of video popularity against its age to determine how relevant it is to recommend to users (Covington et al., 2016). These patterns suggest that the API may similarly sample from topic-wide empirical distributions in terms of over-time interest.
Overall, we do not find evidence of ceiling effects driving the API randomization. Instead, our results suggest that the YouTube API samples videos from empirical distributions, returning results based on the relative density of topical interest and even forcing zero videos to be returned when this relative density is adequately low. However, it is unclear how this interest is computed (e.g., if it is the volume of videos uploaded or something else).
4.3. Attrition Analysis
@@
Figure 3. Transition probabilities for presence (P) or absence (A) of videos in a second-order Markov chain. Example interpretation: Column $A\rightarrow A$ and row $P$ show the probability of P given that the two previous states are A.
@@
5. Factors Behind Return Likelihood
@@
Our next analysis focuses on whether the YouTube API is more likely to return videos with certain features at higher rates. We treat this as an explainability problem where we count the number of times each video is returned in our API calls. We then determine the features that are most predictive of this frequency. Our candidate predictors are topic, video duration, video definition, number of views, comments, and likes at the video level; channel views, subscribers, and number of videos uploaded at the channel level. All continuous features are log-transformed to normalize their distributions.
We deploy a gradient boosting model using LightGBM (Ke et al., 2017) with a Poisson regression objective, 100 estimators, and a learning rate of 0.1. The data are split into 80-20 train-test sets ($R^{2}=0.19)$. We examine the role of each feature from the predictions on the test set using SHapley Additive exPlanations (SHAP) (Lundberg and Lee, 2017), shown in Figure 4. Features are shown in descending order of importance. We also perform robustness checks using regression models with the same variables in Appendix D, which show directionally consistent results.
Figure 4. Beeswarm plot of SHAP values per feature.
Starting with video metadata, we see that the number of views tends to positively predict return frequency and is the most important feature. In terms of duration, which is also among the most important features, shorter videos tend to have more returns. There is also a slightly positive effect of likes, while the number of comments is not as important and video quality plays no role. Overall, this suggests that the API tends to return shorter, more popular videos.
For channel features, returns are more likely from older channels with fewer videos. There is a slightly negative trend in terms of the number of channel subscriptions, while the role of accumulated number of channel views is more mixed. Overall, channel popularity is not as important, though other channel features like age and activity may play a role.
Interestingly, the Higgs Boson and Brexit topics are highly important and positive features for return frequency; the Grammy’s topic also trends towards positive effects. The Capitol riots, BLM, and World Cup topics trend toward negative effects, though the latter two are not as important.
@@
Although the directionally grouped topics look unrelated at first, a deeper look reveals that they are linked by one factor: size. We determine this using metadata returned with every query we send, which contain a value called pageInfo.totalResults. This reflects the total number of results in the result set, i.e., the total number of videos that match our query, with a maximum potential value of 1,000,000. We obtain the minimum, maximum, mean, and modal values from this field across every hour and collection run that we query per topic. As seen in Table 3, the three topics whose videos have higher appearance frequencies are also the smallest (and the only ones without a modal value of 1M, which is the maximum). The results indicate that queries with a smaller video pool to draw from may return more consistent results (and may explain why Higgs, which is by far the smallest topic, is also the most consistent in Figure 1), though this remains to be experimentally verified.
@@
Table 3. Potential video pool size per topic.
@@
Notably, despite the modal number of videos returned across all collection hours for all topics being 0, the modal value for the available pool is much higher (and the maximum of 1M for 3/6 topics). Since it is unlikely that 1M videos are consistently uploaded every hour for any of these topics, this suggests that time constraints in queries do not affect the total pool of available videos in the API’s results set. Rather, they may only filter the response after results are returned.
@@
Comparing Tables 1 and  3, it is also striking that the number of videos returned is much closer across topics than their respective topic sizes suggest, which is consistent with our distribution density explanation. This becomes even more apparent when scrutinizing the y-axis on Figure 2: The most-populated peaks are recorded for topics where the rest of the time-series is relatively inactive (e.g., Capitol Riot, Grammys), while topics that are more active throughout (e.g., World Cup, which is an ongoing tournament rather than a one-off event) record peaks at lower absolute values, most likely due to the number of videos to be returned being fixed and the actual videos being drawn from an empirical distribution.
6. Discussion